40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Feb 7 09:00:26 2022, Paco, Update, General, Scheduled power outage recovery 
    Reply  Mon Feb 7 15:32:48 2022, Yehonathan, Update, General, Y End laser controller 
    Reply  Mon Feb 7 16:53:02 2022, Koji, Update, General, Scheduled power outage recovery PXL_20220208_001646282.jpgPXL_20220208_001657871.jpg
       Reply  Wed Feb 9 11:56:24 2022, Anchal, Update, General, Bringing back CDS Screenshot_2022-02-09_12-11-33.png
          Reply  Wed Feb 9 13:55:05 2022, Koji, Update, General, Bringing back CDS 
       Reply  Wed Feb 9 16:43:35 2022, Paco, Update, General, Scheduled power outage recovery - Locking mode cleaner(s) 
          Reply  Thu Feb 10 15:41:00 2022, Anchal, Update, General, Scheduled power outage recovery - Locking mode cleaner(s) 
             Reply  Thu Feb 10 17:57:48 2022, Anchal, Update, General, Scheduled power outage recovery - Locking mode cleaner(s) PXL_20220211_021509819.jpg
                Reply  Thu Feb 10 19:03:23 2022, Koji, Update, General, Scheduled power outage recovery - Locking mode cleaner(s) 
                   Reply  Thu Feb 10 19:46:37 2022, Koji, Update, General, Scheduled power outage recovery - Locking mode cleaner(s) 
                      Reply  Thu Feb 10 21:10:43 2022, Koji, Update, General, Video Mux setting reset Screenshot_2022-02-10_21-11-21.pngScreenshot_2022-02-10_21-11-54.png
             Reply  Fri Feb 11 16:09:11 2022, Anchal, Update, General, Scheduled power outage recovery - Input power increased 
                Reply  Mon Feb 14 18:31:50 2022, Paco, Update, General, Scheduled power outage recovery - IMC recovery progress 
                   Reply  Tue Feb 15 19:32:50 2022, Koji, Update, General, Scheduled power outage recovery - IMC recovery progress PXL_20220216_001731377.jpgScreen_Shot_2022-02-15_at_16.18.16.pngPXL_20220216_001727465.jpgScreen_Shot_2022-02-15_at_16.22.16.pngPXL_20220216_002229572.jpg
                      Reply  Tue Feb 15 19:40:02 2022, Koji, Update, General, IMC locking 
                         Reply  Wed Feb 16 15:19:41 2022, Anchal, Update, General, Reconfigured MC reflection path for low power 
                            Reply  Wed Feb 23 15:08:57 2022, Anchal, Update, General, Removed extra beamsplitter in MC WFS path 
                               Reply  Thu Feb 24 14:32:57 2022, Anchal, Update, General, MC RFPD DCMON channel got stuck to 0 
                                  Reply  Thu Feb 24 19:26:32 2022, Anchal, Update, General, IMC Locking 
                                     Reply  Sun Feb 27 00:37:00 2022, Koji, Update, General, IMC Locking Recovery PXL_20220226_093809056.jpgPXL_20220226_093854857.jpgPXL_20220226_100859871.jpgScreenshot_2022-02-26_01-56-31.pngScreenshot_2022-02-26_01-56-47.png
                                        Reply  Sun Feb 27 01:12:46 2022, Koji, Update, General, IMC manual alignment procedure PXL_20220226_100859871.jpg
    Reply  Fri Feb 11 11:17:00 2022, Anchal, Update, General, Scheduled power outage recovery 
       Reply  Mon Feb 14 21:03:25 2022, Koji, Update, General, Scheduled power outage recovery PXL_20220215_025325118.jpg
Message ID: 16685     Entry time: Sun Feb 27 00:37:00 2022     In reply to: 16679     Reply to this: 16686
Author: Koji 
Type: Update 
Category: General 
Subject: IMC Locking Recovery 

Summary:

- IMC was locked.
- Some alignment change in the output optics.
- The WFS servos working fine now.
- You need to follow the proper alignment procedure to recover the good alignment condition.

Locking:
- Basically followed the previous procedure 40m/16673.
- The autolocker was turned off. Used MC2 and MC3 for the alignment.
- Once I hit the low order modes, increased the IN1 gain to acquire the lock. This helped me to bring the alignment to TEM00
- Found the MC2 spot was way too off in pitch and yaw.
- Moved MC1/2/3 to bring the MC2 spot around the center of the mirror.
- Found a reasonably good visibility (<90%) at a MC2 spot. Decided this to be the reference (at least for now)

SP Table Alignment Work
- Went to the SP table and aligned the WFS1/2 spots.
- I saw no spot on the camera. Found that the beam for the camera was way too weak and a PO mirror was useless to bring the spot on the CCD.
- So, instead, I decided to catch an AR reflection of the 90% mirror. (See Attachment 1)
- This made the CCD vulnerable to the stronger incident beam to the IMC. Work on the CCD path before increasing the incident power.

MC2 end table alignment work
- I knew that the focusing lens there and the end QPD had inconsistent alignment.
- The true MC2 spot needs to be optimized with A2L (and noise analysis / transmitted beam power analysis / etc)
- So, just aligned the QPD spot using today's beam as the temporary target of the MC alignment. (See Attachment 2)

Resulting CCD image on the quad display (Attachment 3)

WFS Servo
- To activate the WFS with the low transmitted power, the trigger threshold was reduced from 5000 to 500. (See Attachment 4)
- WFS offset was reset with /opt/rtcds/caltech/c1/scripts/MC/WFS/WFS_RF_offsets
- Resulting working state looks like Attachment 5

Attachment 1: PXL_20220226_093809056.jpg  3.877 MB  | Hide | Hide all
PXL_20220226_093809056.jpg
Attachment 2: PXL_20220226_093854857.jpg  3.021 MB  | Hide | Hide all
PXL_20220226_093854857.jpg
Attachment 3: PXL_20220226_100859871.jpg  1.844 MB  | Hide | Hide all
PXL_20220226_100859871.jpg
Attachment 4: Screenshot_2022-02-26_01-56-31.png  86 kB  | Hide | Hide all
Screenshot_2022-02-26_01-56-31.png
Attachment 5: Screenshot_2022-02-26_01-56-47.png  27 kB  | Hide | Hide all
Screenshot_2022-02-26_01-56-47.png
ELOG V3.1.3-